Staff:

LBTO: Steve, Andrew, Juan CArlos

Observer: LBTB ( remote control room)

Forecast.

Screen Shot 2024-04-27 at 4.07.43 PM.png

Screen Shot 2024-04-27 at 4.08.06 PM.png

Details.

04:00 PM

Power on

Ocam Not live.

reconfigure.

masterdiagnostic: OPTLOOPBUF_timeout
  • powercycle bcu2k+reconfigure.
  • OK
AO log:

https://docs.google.com/document/d/1aJzs4qs_ezFE_-yZzo5kJqo2bUrObKkOzP3jQV4kG54/edit?usp=sharing

change esm autogain file:

lrwxrwxrwx. 1 AOeng AOeng 26 Mar 20 21:58 autogain.conf~ -> autogain.conf.ace.20230101
-rw-rw-r--. 1 AOeng AOeng 2565 Mar 25 05:18 daytime.conf
-rw-rw-r--. 1 AOeng AOeng 7072 Mar 25 18:50 wfsarb.conf
lrwxrwxrwx. 1 AOeng AOeng 25 Apr 27 23:21 autogain.conf -> utogain.conf.esm.20230203

close at sunset due to weather.

Starting to observe seeing limited.

02 UT

OPen

seeing above 1.5" variable

Dome closed at ~02:30 due to high humidty

No observing AO

-- JuanCarlosGuerra - 27 Apr 2024
Topic revision: r4 - 29 Apr 2024, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback